OpenStudy (anonymous):

Given that f(x) = x² + 5x, evaluate f(-2). -6

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

f(-2) = (-2)² + 5(-2) = 4 - 10 = -6

OpenStudy (anonymous):

f(x)=x^2+5x f(-2)=(-2)^2+5(-2)=-6